pls explain keplers laws

Grade:9

2 Answers

Vikas TU
13 years ago

Kepler's laws are:

  1. Laws of orbit - Each planet moves around the sun in an elliptical orbit with the sun at one of the foci.
  2. Laws of Areas - The line joining the sun and a planet sweeps out equal areas in equal intervals of time. A planet takes the same time to travel from A to B as from C to D  as shown in figure if the shaded areas are equal. Naturally the planet has to move faster C to D. The law of areas is identical with the law of conservation of angular momentum. Area velocity = area swept/time
  3. Law of Periods - The square of the time for the planet to complete a revolution about the sun is proportional to the cube of semimajor axis of the elliptical orbit.                                                                                                 i.e Centripetal force = Gravitational force

           mv2/R = GMm/R2 = GM/R = v2

Or simply u can understand -

  1. The orbit of every planet is an ellipse with the Sun at one of the two foci.
  2. A line joining a planet and the Sun sweeps out equal areas during equal intervals of time.[1]
  3. The square of the orbital period of a planet is directly proportional to the cubeof the semi-major axis of its orbit.

vikas askiitian expert
13 years ago

dear shivam ,

these all laws are applicable for planatery motion ...

keplers law =>

1) whenever a particle is moving in an elleptical path about a point (called fucus) due to gravitational forces

  then always square of time period is proportional to cube length of semi major axis ...

 

this law is useful to compare the time periods of two different planets which are rotating about sun...

 

2)areal velocity is always constant ....